OpenStudy (anonymous):

Is 3x+y^2=4 a graph of a function??? Please help!!!

Parth (parthkohli):

Yeah it is!

Parth (parthkohli):

\( \color{Black}{\Rightarrow y^2 + 3x - 4 = 0}\)

Parth (parthkohli):

Or am I forgetting my Math?

Parth (parthkohli):

\( \color{Black}{\Rightarrow y^2 = 4 - 3x}\) \( \color{Black}{\Rightarrow y = \sqrt{4 - 3x}}\)

OpenStudy (anonymous):

Solve for y. A plot of the result is attached.

OpenStudy (anonymous):

F(x) = 3x+1,x<=0 x^2,0<x<4,evaluate f(3) X-5,x>=4

OpenStudy (anonymous):

So overall it is a function